OpenStudy (anonymous):

Differentiate the following function with respect to x (1) In2x

sam (.sam.):

1/x

OpenStudy (anonymous):

\[\frac{d}{dx}\ln(ax)=\frac{1}{x}\] for any constant a

OpenStudy (anonymous):

I'm not sure of the rules of In... Are they the same or different than e?

OpenStudy (anonymous):

that is because \[\ln(ax)=\ln(a)+\ln(x)\] the derivative of \[\ln(a)\] is zero, because it is a constant, and the derivative of \[\ln(x)\] is \[\frac{1}{x}\]

OpenStudy (anonymous):

\[ln(x)=\log_e(x)\]

OpenStudy (anonymous):

Hm Ok.. so, what would In(2x-1) be?

OpenStudy (anonymous):

\[\frac{1}{2x-1}\times \frac{d}{dx}[2x-1]\] by the chain rule, or \[\frac{2}{2x-1}\]

sam (.sam.):

2/(2x-1)

OpenStudy (anonymous):

in general the derivative of \[\ln(g(x))\] is \[\frac{g'(x)}{g(x)}\]

OpenStudy (anonymous):

you will use this repeatedly, so it is good to get it under your belt now

OpenStudy (anonymous):

\[what~ does~ ~g'(x) ~mean?\]

OpenStudy (anonymous):

(lnu)' = u'/u (ln 2x)' = 2/2x = 1/x In(2x-1)' = 2/(2x -1)

OpenStudy (anonymous):

note that this also says that since \[\frac{d}{dx}\ln(g(x))=\frac{g'(x)}{g(x)}\]this also means we can solve for g' to get \[g(x)\times \frac{d}{dx}\ln(g(x))=g'(x)\] a good way to find the derivative sometimes

OpenStudy (anonymous):

g(x) = u

OpenStudy (anonymous):

\[g'\] means the derivative of g

OpenStudy (anonymous):

Hmm. Ok

OpenStudy (anonymous):

I've done some more on my own, and have managed, but what's\[ In {1 \over x}?\]

sam (.sam.):

-1/x

OpenStudy (anonymous):

Yes, but why?

sam (.sam.):

\[y=\ln\frac{1}{x}\] \[\frac{dy}{dx}=-\frac{1}{(\frac{1}{x})}(x ^{-2})\] \[\frac{dy}{dx}=-x(x ^{-2})\] \[\frac{dy}{dx}=-\frac{x}{x ^{2}}\] \[\frac{dy}{dx}=-\frac{1}{x}\]

OpenStudy (anonymous):

Oh, That makes soo much more sense! Thanks so much!